OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2

OpenStudy (anonymous):

Use point-slope formula:\[y-y _{0}=m(x-x _{0})\] So, y - (-4) = 1/2 (x - 1) If you want it in the form y=mx+b, then just solve solve y. y + 4 = 1/2x - 1/2 y = 1/2x - 9/2

OpenStudy (anonymous):

To check your answer, insert the given (x,y) values are confirm the equation.

OpenStudy (anonymous):

thank you